Course Management: A complete asynchronous environment architecture in education

Mehmet Sen, Northeast Parallel Architectures Center, Syracuse University, NY, USA. E-mail: msen7@npac.syr.edu

Ozgur Balsoy, Northeast Parallel Architectures Center, Syracuse University, USA. E-mail: balsoy@npac.syr.edu

Geoffrey C. Fox, Northeast Parallel Architectures Center, Syracuse University, USA. E-mail: gcf@npac.syr.edu

Nancy J. McCracken, Northeast Parallel Architectures Center, Syracuse University, USA. E-mail: njm@npac.syr.edu

July 4, 1999

Abstract

Scheduling computation tasks on processors is the key issue for highperformance computing. Although a large number of scheduling heuristics have been presented in the literature, most of them target only homogeneous resources. The existing algorithms for heterogeneous domains are not generally efficient because of their high complexity and/or the quality of the results. We present two low-complexity efficient heuristics, the Heterogeneous Earliest-Finish-Time (HEFT) Algorithm and the Critical-Path-on-a-Processor (CPOP) Algorithm for scheduling directed acyclic weighted task graphs (DAGs) on a bounded number of heterogeneous processors. We compared the performances of these algorithms against three previously proposed heuristics. The comparison study showed that our algorithms outperform previous approaches in terms of performance (schedule length ratio and speedup) and cost (time-complexity).